Could you please help me to solve this exercise, I m very grateful if you can leave the formula of calculation in details so that I can understand how to solve myself for other similar exercises
Many thanks
CAB Company manufactures two different products: X and Y. The following information is known: Quantity made Direct labor (per unit) Number of components (per unit) Product X 10,000 2 minutes Product Y 2,000 4 minutes 25 Total fixed manufacturing overhead are $ 280,000 of which $ 224.000 ar incurred in order to perform activities caused by the number of components. Required: (1) Compute the manufacturing fixed cost per unit of Product X an Product Y using the traditional costing systems approach. (2) Compute the same cost figures using ABC
Answer #1

Answer (1):

In Traditional costing approach, manufacturing fixed cost will be allocated based on direct labor hour.

First let us calculated rate to be applied = Total fixed manufacturing overhead / total direct labor in minutes

= $280,000 / ( 2 * 10,000 + 4 * 2000)

= $280,000 / 28,000 minutes

= $10 per minute

As such:

Manufacturing fixed cost per unit of Product X = $10 * 2 minutes = $20

Manufacturing fixed cost per unit of Product Y = $10 * 4 minutes = $40

Manufacturing fixed cost per unit of Product X = $20

Manufacturing fixed cost per unit of Product Y = $40

Answer (2)

ABC approach:

Since there cause and effect established that:

Manufacturing fixed cost of $224,000 is incurred in order to perform activities caused by number of components.

As such we will use number of components as cost driver to allocate $224,000

Total components = 10,000 * 5 + 2,000 * 25 = 100,000

Cost driver rate per component = $224,000 / 100,000 = $2.24 per component

Balance fixed manufacturing cost = $280,000 - $224,000 = $56,000

Balance fixed manufacturing cost will be allocated based on direct labor hour

Rate to be applied = $56,000 / ( 2 * 10,000 + 4 * 2000) = $2 Per direct labor minute

Hence:

Manufacturing fixed cost per unit of Product X = $2.24* 5 + $2 * 2 = $15.20

Manufacturing fixed cost per unit of Product Y = $2.24* 25 + $2 * 4 = $64

Manufacturing fixed cost per unit of Product X = $15.20

Manufacturing fixed cost per unit of Product Y = $64